# Uncovering Grounding IDs: How External Cues Shape Multimodal Binding

### v3

@@ A.1 Visual Traversal and Positional Inductive Bias
As a qualitative example, Figure 10 illustrates cross-attention heatmaps for individual objects. Each panel corresponds to a specific shape–color pair, showing how the model localizes and attends to the correct region during decoding. This highlights the row-by-row traversal pattern on a per-object basis.
@@
A.2 Logit Lens Analysis and Evidence for Propagation
As shown in Figure 12, the manipulated images (left) produce sharper symbol activations that extend beyond their immediate positions, spreading across all patches within the corresponding row. This demonstrates how symbolic structure strengthens local grounding while also enabling coherent row-level propagation. By comparison, the baseline condition (right) lacks this structured diffusion, resulting in more diffuse and inconsistent attention patterns. Together, these findings highlight the role of explicit symbols in amplifying the model’s evidence propagation and improving cross-modal grounding.
A.3 Intra-Visual Attention Patterns
@@
where $R$ is the number of rows, $\text{Sim}_{\mathrm{intra}}(r)$ is the mean cosine similarity between patches within row $r$, and $\text{Sim}_{\mathrm{inter}}(r,r^{\prime})$ is the average similarity between patches across rows $r$ and $r^{\prime}$. A larger $\mathrm{ICG}$ score indicates stronger row-wise clustering and hence a clearer positional alignment in the image representation. Figure 13 plots the ICG score across layers for both the baseline and the structured setting. We observe that the structured condition (symbol-augmented inputs) consistently achieves higher ICG values than the baseline, indicating that structural cues enhance row-level grouping.
@@
A.4 Evaluation on Synthetic Variants
@@
Grid 4×4 Numbered 0.33 0.40 0.36 0.22
Grid 4×4 Numbered (no lines) 0.33 0.35 0.34 0.20
@@
A.5 Implementation Details and Qualitative Example
@@
Output format (print exactly this structure, no extra text):
@@
Structured Prompt Output:
Row &: purple moon, purple star, blue moon
@@
Row @: blue circle, purple circle
@@
Baseline Prompt Output:
purple moon, purple star, purple diamond, yellow heart,
@@
A.6 Evaluation Benchmarks for Object Hallucination
@@
A.6.1 CHAIR: Caption Hallucination Assessment with Image Relevance
@@
A.6.2 POPE: Probing Object-Presence Hallucination
@@
A.6.3 Evaluation on POPE Benchmark
In addition to CHAIR, we evaluate our structured scaffolding method using the POPE benchmark, which probes object hallucination through structured binary questions. As described in Appendix A.6.2, POPE defines three subsets—Random, Adversarial, and Popular—each targeting different hallucination biases such as language priors, scene co-occurrence, and object frequency.
@@ A.6.3 Evaluation on POPE Benchmark
A.7 Hyperparameters for Hallucination-Mitigation Methods
@@
A.8 Activation swapping (interchange intervention)
@@
A.8.1 Disjoint-symbol swap (no-visual-cue prompt)
@@
A.9 Extended Experiments on Alternative Models
Experiments in Sections 3 and 4 were originally conducted on Qwen2.5-VL (7B). Here, we present additional experiments on models with different sizes and architectures. Figures 15 and 16 show within- and cross-modality attention analyses for InternVL3.5 (8B) and Qwen2.5-VL (3B), which exhibit the same attention progression as Qwen2.5-VL (7B) (Figure 2).
@@ A.9 Extended Experiments on Alternative Models
We also repeated the activation-swap experiment from Section 4 on InternVL3.5 (8B) and Qwen2.5-VL (3B). The models exhibit the same qualitative behavior as Qwen2.5-VL (7B). For InternVL3.5, swap accuracy is 0.75 for shape and 0.82 for color, well above the random chance level of 0.25 for shape and 0.125 for color. For Qwen2.5-VL (3B), swap accuracy is 0.93 for shape and 0.73 for color, indicating a consistent dependence on Grounding IDs. Figure 17 shows the corresponding log-prob diagrams, which follow the same pattern observed in Qwen2.5-VL (7B) (Figure 4(b)).

### v4

@@ Appendix B Visual Traversal and Positional Inductive Bias
As a qualitative example, Fig. 10 illustrates cross-attention heatmaps for individual objects. Each panel corresponds to a specific shape–color pair, showing how the model localizes and attends to the correct region during decoding. This highlights the row-by-row traversal pattern on a per-object basis.
@@
Appendix C Intra-Visual Attention Patterns
@@
where $R$ is the number of rows, $\text{Sim}_{\mathrm{intra}}(r)$ is the mean cosine similarity between patches within row $r$, and $\text{Sim}_{\mathrm{inter}}(r,r^{\prime})$ is the average similarity between patches across rows $r$ and $r^{\prime}$. A larger $\mathrm{ICG}$ score indicates stronger row-wise clustering and hence a clearer positional alignment in the image representation. Fig. 12 plots the ICG score across layers for both the baseline and the structured setting. We observe that the structured condition (symbol-augmented inputs) consistently achieves higher ICG values than the baseline, indicating that structural cues enhance row-level grouping.
@@
Appendix D Logit Lens Analysis and Evidence for Propagation
As shown in Fig. 13, the manipulated images (left) produce sharper symbol activations that extend beyond their immediate positions, spreading across all patches within the corresponding row. This demonstrates how symbolic structure strengthens local grounding while also enabling coherent row-level propagation. By comparison, the baseline condition (right) lacks this structured diffusion, resulting in more diffuse and inconsistent attention patterns. Together, these findings highlight the role of explicit symbols in amplifying the model’s evidence propagation and improving cross-modal grounding.
Appendix E Evaluation on Synthetic Variants
@@
Grid $\times4$ Numbered 0.33 0.40 0.36 0.22
Grid $\times4$ Numbered (no lines) 0.33 0.35 0.34 0.20
@@
Appendix F Implementation Details and Qualitative Example
@@
Structured Prompt Output: Row &: purple moon, purple star, blue moon
@@
Row : blue circle, purple circle
@@
Baseline Prompt Output: purple moon, purple star, purple diamond, yellow heart,
@@
Appendix G Evaluation Benchmarks for Object Hallucination
@@
Appendix H CHAIR: Caption Hallucination Assessment with Image Relevance
@@
Appendix I POPE: Probing Object-Presence Hallucination
@@
Appendix J Evaluation on POPE Benchmark
In addition to CHAIR, we evaluate our structured scaffolding method using the POPE benchmark, which probes object hallucination through structured binary questions. As described in Appendix I, POPE defines three subsets—Random, Adversarial, and Popular—each targeting different hallucination biases such as language priors, scene co-occurrence, and object frequency.
@@
Appendix K Hyperparameters for Hallucination-Mitigation Methods
@@
Appendix L Activation swapping (interchange intervention)
@@
Appendix M Disjoint-symbol swap (no-visual-cue prompt)
@@
Appendix N Extended Experiments on Alternative Models
Experiments in Sections 3 and 4 were originally conducted on Qwen2.5-VL (7B). Here, we present additional experiments on models with different sizes and architectures. Figures 15 and 16 show within- and cross-modality attention analyses for InternVL3.5 (8B) and Qwen2.5-VL (3B), which exhibit the same attention progression as Qwen2.5-VL (7B) (Fig. 2).
@@
We also repeated the activation-swap experiment from Section 4 on InternVL3.5 (8B) and Qwen2.5-VL (3B). The models exhibit the same qualitative behavior as Qwen2.5-VL (7B). For InternVL3.5, swap accuracy is 0.75 for shape and 0.82 for color, well above the random chance level of 0.25 for shape and 0.125 for color. For Qwen2.5-VL (3B), swap accuracy is 0.93 for shape and 0.73 for color, indicating a consistent dependence on Grounding IDs. Fig. 17 shows the corresponding log-prob diagrams, which follow the same pattern observed in Qwen2.5-VL (7B) (Fig. 4(b)).